An electric drill draws 4.5 A rms at 150 V rms. If the current lags the voltage by 45°, what is the drill's power consumption?
W
Answer #1

V = 150 Volts

I = 4.5 A

theta = 45

power consumption = P =V*I*cos(theta) = 150*4.5*cos(45) =477.29 Watts